OpenStudy (mendicant_bias):

Is there some sort of convention of written word about ratios, like "Show that the ratio of x to y...", does that mean x/y, or y/x generally? I've ran into a problem where they bring up a ratio, but it's the second case, which seems a little strange.

OpenStudy (texaschic101):

ratio of x to y = x/y or x:y

OpenStudy (texaschic101):

ratio of y to x = y/x or y:x

OpenStudy (mendicant_bias):

"Show that the ratio of dh/dt to the rate of rainfall is equal to the ratio between the area of the tank’s opening to the area of the water surface." They then compare it like \[\frac{ k }{ \frac{ dh }{ dt } }\] (Where k is the rate of rainfall) which seems a little unexpected, I thought it would be kind of a, first stated thing is in the numerator, second is in the denominator, deal.

OpenStudy (phi):

I would put dh/dt in the top, and rate of rainfall in the bottom

OpenStudy (mendicant_bias):

Okay, so it's just poor writing. Thanks, guys.